# Reliability in Semantic Segmentation: Can We Use Synthetic Data?

### v1

@@ A.5 Segmentation Models
We list here all models used in our experiments: ANN-R101, ANN-R50, APCNet-R101, APCNet-R50, BiSeNetV1-R50, BiSeNetV2-FCN, CCNet-R101, CCNet-R50, ConvNext, ConvNext-B-In1K, ConvNext-B-In21K, DLV3+ResNet101, DLV3+ResNet18, DLV3+ResNet50, GCNet-R101, GCNet-R50, ICNet-R101, ICNet-R18, ICNet-R50, MobileNetV3, PSPNet-R101, PSPNet-R18, PSPNet-R50, SETR-MLA, SETR-Naive, SETR-PUP, SegFormer-B0, SegFormer-B1, SegFormer-B2, SegFormer-B3, SegFormer-B4, SegFormer-B5, SegFormer-B5-v2, SegFormer-B5-v3, Segmenter, SemFPN-R101, SemFPN-R50, UpperNet-R101, UpperNet-R18, UpperNet-R50.
Appendix B Calibration Details and Additional Results
@@
We elaborate on our strategy for performing per-class calibration to obtain the synthetic results ( ) presented in Figure 6. Utilizing our synthetic data, a temperature scaling (TS) scalar is learned for each class. When calibrating models on shifted domains, we choose the corresponding TS scalar based on model predictions. In the case of calibration results with real-shift data ( ), only one scalar is learned for each model.
@@
In 12(a), we compare per-class TS vs. standard TS with one scalar per model, both applied on our synthetic data. Both strategies enhance calibration results, highlighting the advantage of employing synthetic data for calibration. Per-class TS demonstrates superiority for more robust models (right part of the plots), while its performance is weaker for less robust ones (left part of the plots).
@@
In 12(b), we compare Cityscapes vs. our synthetic data; in this experiment we adopt the standard TS with one scalar per model. The results obtained from Cityscapes are clearly inferior to those achieved using our synthetic data, demonstrating the limitations when relying solely on in-domain data for calibration in shifted domains.
@@
(a) Perclass Temperature Scaling ( ) vs. One Temperature Scaling ( ). The Figure has the same structure as of Figure 6 and the bars show relative ECE improvements. We compare the two strategies for performing calibration using synthetic data; both enhance calibration in shifted domains.
@@
(b) Cityscapes ( ) vs. our synthetic data with one TS ( ). The Figure has the same structure as of Figure 6 and the bars show relative ECE improvements. Our synthetic data is superior to Cityscapes in improving calibration in shifted domains.
@@
Appendix C Qualitative Examples
We show more qualitative examples for synthetic covariate shifts in Figure 14 and synthetic OOD objects in Figure 15.
Synthetic Images Semantic-FPN MobileNetV3 SegF-B5 Synthetic Images Semantic-FPN MobileNetV3 SegF-B5
@@ Appendix C Qualitative Examples
Figure 15: Qualitative results. Confidence maps are visualized for the four exemplified models on real data (first row) and synthetic inpainted data (second and third rows). Hotter colors correspond to higher OOD likelihood. Ideally, results should exhibit hot colors in OOD areas and cold colors everywhere else. We observe a strong correlation in model reactions to real and synthetic OOD regions, particularly for more recent and robust models.

### v2

@@ 0.A.5 Segmentation Models
We list here all models used in our experiments: ANN-R101, ANN-R50, APCNet-R101, APCNet-R50, BiSeNetV1-R50, BiSeNetV2-FCN, CCNet-R101, CCNet-R50, ConvNext, ConvNext-B-In1K, ConvNext-B-In21K, DLV3+ResNet101, DLV3+ResNet18, DLV3+ResNet50, GCNet-R101, GCNet-R50, ICNet-R101, ICNet-R18, ICNet-R50, MobileNetV3, PSPNet-R101, PSPNet-R18, PSPNet-R50, SETR-MLA, SETR-Naive, SETR-PUP, SegFormer-B0, SegFormer-B1, SegFormer-B2, SegFormer-B3, SegFormer-B4,
SegFormer-B5, SegFormer-B5-v2, SegFormer-B5-v3, Segmenter, SemFPN-R101,
SemFPN-R50, UpperNet-R101, UpperNet-R18, UpperNet-R50.
Appendix 0.B Calibration Details and Additional Results
@@
We elaborate on our strategy for performing per-class calibration to obtain the synthetic results ( ) presented in Fig. 6. Utilizing our synthetic data, a temperature scaling (TS) scalar is learned for each class. When calibrating models on shifted domains, we choose the corresponding TS scalar based on model predictions. In the case of calibration results with real-shift data ( ), only one scalar is learned for each model.
@@
In Fig. 14(a), we compare per-class TS vs. standard TS with one scalar per model, both applied on our synthetic data. Both strategies enhance calibration results, highlighting the advantage of employing synthetic data for calibration. Per-class TS demonstrates superiority for more robust models (right part of the plots), while its performance is weaker for less robust ones (left part of the plots).
@@
In Fig. 14(b), we compare Cityscapes vs. our synthetic data; in this experiment we adopt the standard TS with one scalar per model. The results obtained from Cityscapes are clearly inferior to those achieved using our synthetic data, demonstrating the limitations when relying solely on in-domain data for calibration in shifted domains.
Generalization. We ask ChatGPT the generic question like “give me different cities/weathers that would be representative of the whole world?”. We then use all ChatGPT’s answers as prompts to generate OOD data, referred to as “all-domains”. For confidence calibration, using “all-domains” shows comparable results as using domain-specific data with manual prompts (see Tab. 4), e.g. $100\%$ ECE improvement for rain, $78.6\%$ for india (cf. Fig.6). Interestingly, using “all-domains-but-rain” (no rain prompts), we also obtain similar improvement for rain. Detailed results are reported in Tab. 3. That experiment hints at the generalization potential of our framework. We list here all the domains we used to form the “all-domains” calibration set: Beijing, Cairo, clouds, Dubai, fall, fog, hurricane, India, Istanbul, Johannesburg, lightning, London, Moscow, Mumbai, night, Paris, rain, sandstorm, snow, spring, summer, sun, Sydney, Tokyo, tornado, Toronto, wind, winter. This calibration set is comprised of 64 images per shift.
Prompt India Fog Rain Snow Night
domain-specific 72.5 92.5 100 95 90
all domains 78.6 100 100 100 100
all domains but-rain - - 100 - -
Table 3: Generalization. We experiment the effects of prompting and showcase the generalization potentials of our framework. Here ECE improvements are reported (cf. Fig.6).
@@
(a) Per class Temperature Scaling ( ) vs. One Temperature Scaling ( ). The Figure has the same structure as of Fig. 6 and the bars show relative ECE improvements. We compare the two strategies for performing calibration using synthetic data; both enhance calibration in shifted domains.
@@
(b) Cityscapes ( ) vs. our synthetic data with one TS ( ). The Figure has the same structure as of Fig. 6 and the bars show relative ECE improvements. Our synthetic data is superior to Cityscapes in improving calibration in shifted domains.
@@
Appendix 0.C Direct Data Correlation
We provide in Tab. 4 the FID scores($\downarrow$) for direct correlation between synthetic and real data. Our zero-shot approach outperforms all. Layout differences between Cityscapes and ACDC cause high FIDs – one critical limitation of FID on structured data. As our work focuses on performance testing, we prioritize metrics like mIoU and FPR95, aligned with recent works.
OOD expertise? OOD data? Night Rain Snow Fog
Cityscapes (no aug.) no no 236.5 188.3 194.8 184.8
Hendrycks’ aug. [11] required no - - 210.7 191.1
GAN-based TSIT [13] no required 254.2 223.2 225.6 -
Physics-based Fog Sim. [32] required no - - - 182.8
Ours w/ SD1.5 no no 180.2 177.5 164.5 163.3
Table 4: Direct Data Correlation. Our pipeline achieves better FID scores while does not require any OOD knowledge.
Reflected in the FID study, in-domain data (CS) serves as a strong baseline for measuring OOD performance [14, 23]. Hendrycks’ augmentations are unrealistic and inadequate for assessing real OOD performance, as previously revealed in [36]. Using untouched CS data could be even better; for example, in terms of PCC score for ‘fog’ (cf. Fig.3 and Tab.1), Hendrycks’s data obtains 0.70 as compared to 0.78 of CS data and 0.89 of our SDXL data. Our work, consistent with [36], demonstrates that simple augmentations are insufficient for OOD testing, advocating for more realistic testing data. While showing the zero-shot advantage over standard domain-specific augmentations, we acknowledge that a well crafted augmentation approach like physics-based fog augmentation [32] can obtain very good results, potentially better than ours if improved. Of note, we do not aim to obtain best results in all cases; we advocate for a generic zero-shot testing framework as the starting reference in arbitrary domains.
Appendix 0.D Class-wise analysis.
We conduct a class-wise analysis and report the PCC scores per class in Tab. 5. Interestingly, we notice that ‘bicycle’ and ‘bus’ have the least correlation for india and night, respectively, which actually corresponds to the low occurrences of those classes in such conditions. High correlation indicates that either the corresponding classes are easy (‘building’, ‘person’, ‘car’, or ‘truck’) or those classes are difficult (’sign’ or ’pole’) and make segmenters struggle either on real or synthetic data. We conjecture that using synthetic data may provide us with hints about the inherent bias of the pretrained models.
Class Fog Night Rain Snow India
road 0.54 0.69 0.57 0.60 0.74
sidewalk 0.62 0.62 0.68 0.64 0.69
building 0.67 0.82 0.76 0.57 0.84
wall 0.03 0.76 0.60 0.71 0.58
fence 0.21 0.66 0.71 0.64 0.38
pole 0.33 0.75 0.60 0.29 0.47
light 0.36 0.53 0.63 0.70 0.30
sign 0.33 0.82 0.47 0.55 0.26
vegetation 0.57 0.71 0.77 0.52 0.00
terrain 0.31 0.55 0.62 0.76 -
sky 0.70 0.44 0.07 0.17 0.30
person 0.56 0.72 0.52 0.69 0.86
rider 0.27 0.55 0.43 0.31 0.16
car 0.73 0.71 0.72 0.78 0.61
truck 0.83 0.67 0.74 0.75 0.72
bus 0.67 0.32 0.57 0.69 0.19
train 0.35 0.65 0.35 0.41 -
motocycle 0.25 0.39 0.37 0.62 0.40
bicycle 0.49 0.64 0.73 0.72 -0.16
Table 5: Class-wise analysis. We provide the PCC scores per class for each shift. The most and least scores are colored.
Appendix 0.E Limitations
We focus our research on the task of semantic segmentation while keeping open the extension possibility to other critical tasks, such as object detection. Our quantitative assessments are confined to existing publicly available datasets. However, our framework is fully zero-shot and can be applied to any domain of interest. On the generative side, our study is restricted to Stable Diffusion and ControlNet due to our resource constraints. Of note, although improvements in this area should enhance the results, similar insights are expected to be achieved, as primarily shown with SDXL results. Another limitation is our primary focus on autonomous driving (AD) data. That is because the field of AD released diverse datasets to stress-test models across various OOD scenarios. Unlike generalist datasets like MS-COCO, they clearly distinguish between domains, enabling the covariate shift studies in this work.
Robustness Assessment. In order to accommodate the in-domain GTs, we make the semantic preservation assumption, similar to the one in domain adaptation. We realize that this common assumption might raise questions when an element like snow is added; nevertheless, the issue is complex and hinges on the annotation policies. The ACDC dataset [33], for instance, uses clean images to inform the annotators of what is behind the snow, and our approach mimics this. But they are also very conservative about ambiguity, with an explicit label. Moving forward, we could take inspiration and try to handle ambiguity.
Appendix 0.F Qualitative Examples
We show more qualitative examples for synthetic covariate shifts in Fig. 15 and synthetic OOD objects in Fig. 16.
Syn. Img Sem.FPN MobileV3 SegF-B5 Syn. Img Sem.FPN MobileV3 SegF-B5
@@ Appendix 0.F Qualitative Examples
Figure 16: Qualitative results. Confidence maps are visualized for the four exemplified models on synthetic inpainted data. Hotter colors correspond to higher OOD likelihood. Ideally, results should exhibit hot colors in OOD areas and cold colors everywhere else.